A supermarket polled 1,000 customers regarding the size of their bill. The results are given in the table below.
Size of Bill Number of Customers
below $20.00 203
$20.00–$39.99 117
$40.00–$59.99 182
$60.00–$79.99 176
$80.00–$99.99 193
$100.00 or above 129
Use probability rules (when appropriate) to find the relative frequency with which a customer's bill is as stated. (Enter your answers as fractions.)

(a) between $20.00 and $79.99




(b) not between $20.00 and $79.99
1
Expert's answer
2012-03-13T07:41:42-0400
The total number of customers is N = 1000.

(a) between $20.00 and $79.99

The number of customers with the bill size between $20.00 and $79.99 is n = 117+182+176 = 475. So, relative frequency is F1 = n/N = 475/1000 = 19/40.
&
(b) not between $20.00 and $79.99

The number of customers with the bill size not between $20.00 and $79.99 is m = N-n = 1000-475 = 525. So, relative frequency is F1 = m/N = 525/1000 = 21/40.